OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 + 2x2 – 24x ?

OpenStudy (lgbasallote):

hint: factor out x x(x^2 + 2x - 24) can you find the zeroes now? or do you need more help?

OpenStudy (klimenkov):

Try to use \(f(x)=x^3+2x^2-24x=x(x^2+2x-24)\).